# Post-detection inference for sequential changepoint localization

### v4

@@ 3 A general framework for nonparametric post-change
Let $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ denote the pre- and post-change classes respectively, $\mathcal{P}_{0}\cap\mathcal{P}_{1}=\emptyset$. Our goal is to construct post-detection confidence sets for the changepoint $T$ by inverting tests for $H_{0,t}$, as formalized in Proposition 2.2.
To motivate the general construction, we begin with the simplest case when both pre- and post-change distributions $F_{0}$ and $F_{1}$ (with densities $f_{0}$ and $f_{1}$) are known. Then, the likelihood of the data $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}$, $X_{T},\cdots,X_{n}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}$ is given by $L_{T}^{n}:=\prod_{i=1}^{T-1}f_{0}(X_{i})\prod_{i=T}^{n}f_{1}(X_{i})$, and the maximum likelihood estimator of $T$ [10] can be written as $\hat{T}_{n}=\operatorname*{arg\,max}_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ In our setting, having data is observed up to a stopping time $\tau$, the analogous estimator is obtained by replacing $n$ with $\tau$ [40, 39]:
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ (3.1)
Now, a prototypical likelihood-ratio-based test statistic is the following: $M_{t}=\displaystyle\max_{1\leq i\leq\tau}{L_{i}^{\tau}}/{L_{t}^{\tau}},\text{ if }t\leq\tau<\infty$ and $M_{t}=-\infty\text{ if }t>\tau\text{ or }\tau=\infty$. It simplifies to
@@
One natural way to extend this test statistic for the general nonparametric setting is to replace the likelihood ratio processes by e-processes [20, 8], which have emerged as a central tool in anytime-valid inference. While e-processes have been used for sequential testing [36, 21], estimation [37], and change detection [28], we employ them for the first time for changepoint localization. For $t\in\mathbb{N}$, we define two different types of e-processes:
@@
Such powered e-processes for $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ (and vice versa) exist if the total variation distance between the two convex hulls of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ is positive (see e.g., Section 3.4 of [22]). The likelihood ratio process is the simplest example of an e-process. There is a substantial literature on constructing e-processes in both parametric (e.g., [36]) and nonparametric settings (e.g., [37]). For readers less familiar with these developments, we refer to Chapters 6 and 7 of the recent textbook [22]. Note that we require all $X_{i}$’s to be independent for both powered (nontrivial) e-processes to be adapted to their respective filtrations.
The final ingredient we need is $\hat{T}$, which is a preliminary point estimate of $T$.  [^4] Here, we again note that our confidence set based on a point estimate $\hat{T}$ can be viewed analogously to the conformal prediction set, which is constructed around a preliminary “point prediction”. One can easily construct a point estimate. For example, let $\{S^{(t)}_{n}\}_{n\geq t}$ be some forward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$). Then, one reasonable estimate is
$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau}S^{(t)}_{\tau}$. (3.3)
This can be viewed as a nonparametric generalization of the maximum likelihood estimator for known pre- and post-change defined in (3.1).
With the above ingredients in place, we construct a test statistic for $H_{0,t}$ by comparing it against a competing hypothesis $H_{0,T}$. Suppose $t<T$. Then under $H_{0,t}$ (which is false), the segment $X_{t},\dots,X_{T-1}$ follows a post-change distribution from $\mathcal{P}_{1}$ (because under $H_{0,t}$, the change occurred at $t$), whereas under $H_{0,T}$ (which is true), it follows a pre-change distribution from $\mathcal{P}_{0}$. Outside this segment $X_{t},\dots,X_{T-1}$, both hypotheses agree on the distribution of the observations, and hence those are non-informative for testing between $H_{0,t}$ and $H_{0,T}$. Therefore, we use a forward e-process $R^{(t)}_{T-1}$ that accumulates evidence in favor of $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ over this segment. Similarly, if $t>T$, we use a backward e-process $S^{(t)}_{T}$ to accumulate evidence for $\mathcal{P}_{0}$ against $\mathcal{P}_{1}$, using the segment $X_{T},\dots,X_{t-1}$ . Since the true $T$ is unknown, we replace it with an estimated change-point $\hat{T}$. Fig. 3.1 provides an illustration of the key idea. This naturally leads us to define:
[…]
Figure 3.1: Pre-change and post-change parts are shown in black and blue, respectively. Outside the time segment $t,\dots,{\hat{T}-1}]$ when $t<\hat{T}$ (or, $\hat{T},\dots,{t-1}]$ when $t>\hat{T}$) both $H_{0,t}$ and $H_{0,\hat{T}}$ agree on the distribution of the observations.
@@
While this bound may be conservative, it is remarkably universal: it holds without any assumptions on the distribution classes $\mathcal{P}_{0}$, $\mathcal{P}_{1}$, or on the detection method $\mathcal{A}$. For constructing a valid confidence set, Proposition 2.2 implies that we require tests of level $\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t)$, whose value is of course not known to us. To circumvent this issue, we first assume that the pre-change distribution $F_{0}$ is known. This is a common assumption in sequential changepoint literature.
@@ 3.1 Confidence set for $T$ with known pre-change distribution
$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}}\right\}.$ (3.5)
@@
Theorem 3.2.
Suppose that the pre-change distribution is known, i.e., $\mathcal{P}_{0}=F_{0}$. Then, for any $\alpha\in(0,1),T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfying $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (3.5), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{1}\in\mathcal{P}_{1}.$
Now we analyze the size of the confidence set (3.5) for known pre-change $F_{0}$ and post-change $F_{1}$, with $r_{t}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)}{N+1}$, $M_{t}$ as defined in (3.2) with $\hat{T}$ as defined in (3.1). We provide a bound on the expected size of the confidence set as $N\to\infty$, where $N$ is the number of simulations used to calculate $r_{t}$. Define, $p_{t}=\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$. Note that $r_{t}\to p_{t}$ as $N\to\infty$. Define $\rho_{0}(s)=\mathbb{E}_{F_{0}}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}$ and $\rho_{1}(s)=\mathbb{E}_{F_{1}}\left(\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right)^{s}$. These two quantities measure closeness between $F_{0}$ and $F_{1}$, and thus the problem hardness. Note that for $s=1/2,$ this coincides with Bhattacharyya distance. At $s=0,1$, $\rho_{0}$ and $\rho_{1}$ take value $1.$ Since for $s\in(0,1)$, $h(x)=x^{s}$ is a strictly concave function and since $F_{1}\neq F_{0}$, $\frac{f_{1}(X)}{f_{0}(X)}$ is not almost everywhere constant, we get a strict inequality using Jensen’s: $\rho_{0}(s)=\mathbb{E}_{F_{0}}\left[\left(\frac{f_{1}(X)}{f_{0}(X)}\right)^{s}\right]<\left(\mathbb{E}_{F_{0}}\left[\frac{f_{1}(X)}{f_{0}(X)}\right]\right)^{s}=1.$ Similarly, for $s\in(0,1),\rho_{1}(s)<1$. We now define $s_{0}=\arg\min_{s\in[0,1]}\rho_{0}(s)$ and $s_{1}=\arg\min_{s\in[0,1]}\rho_{1}(s)$, and note that the preceding facts imply that $0<s_{0}<1$ and $0<s_{1}<1$. Our length bound will depend on the expected detection delay, $\Delta_{F_{0},T,F_{1}}:=\mathbb{E}_{F_{0},T,F_{1}}(|\tau-T|\mid\tau\geq T).$ For typical detection algorithms, this quantity often scales like the logarithm of the average run length divided by the KL divergence between $F_{1},F_{0}$. Finally, we call the stopping time $\tau$ as sensitive if
$\penalty 10000\ \mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]\geq\mathbb{P}_{F_{0},\infty}[\tau\leq t],$ (3.6)
@@
Theorem 3.3.
In the setting described above, for any change detection method with stopping time $\tau$ satisfying and $p_{T}\neq 0$, the size of $\mathcal{C}$} can be bounded as:
$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\cdot\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{F_{0},T,F_{1}}.$ (3.7)
Additionally, if $\tau$ is sensitive in the sense of (3.6), the last term in (3.7) can be improved to $\Psi_{F_{0},T,F_{1}}:=\min\left\{\Delta_{F_{0},T,F_{1}},\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right)\right\}$.
Remark 3.4.
Suppose $F_{1}=N(\mu_{1},\sigma^{2})$ vs $F_{0}=N(\mu_{0},\sigma^{2})$ and we are using a CUSUM detector with the likelihood ratio (which satisfies the assumption) with average run length $A$. Consider the asymptotic regime when $A\to\infty$, and $T=CA$, for some $C>0$. Using the property of the CUSUM detector that $\frac{\tau}{\mathbb{E}_{F_{0},\infty}(\tau)}\stackrel{{\scriptstyle d}}{{\to}}\exp(1)$ as $A\to\infty$, we have $p_{T}=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq CA_{T})\to\exp(-C)$. Also, for the Gaussian distribution, it is easy to verify that $\rho_{0}(s)=\rho_{1}(s)=\exp\left(-s(1-s)K\right)$, where $K=\frac{(\mu_{1}-\mu_{0})^{2}}{2\sigma^{2}}$ and $s_{0}=s_{1}=1/2$. Substituting, we get
@@
To simplify further, consider the difficult regime of small $K$, where $\left(\exp\left(\frac{K}{4}\right)-1\right)\approx K/4$, and $\left(\exp\left(\frac{K}{8}\right)-1\right)\approx K/8$, and take $C=\frac{1}{3}\log(\frac{2}{\alpha})$, so that the above expression is bounded by, ignoring constants, $K^{-1}[\frac{1}{\alpha}+\frac{1}{\alpha^{5/6}}++\frac{1}{\alpha^{1/3}}]$. The fact that the bound grows with $C=T/A$ is empirically verified in Supplementary Section G.1. This behavior is not a limitation of our methodology; rather, it reflects a fundamental difficulty of the problem: $p_{t}$ decays as $C$ increases, and hence, by our duality result (Proposition 2.2), the tests corresponding to any confidence set will have smaller levels $\alpha p_{t}$, leading to reduced power and, consequently, wider confidence sets.
@@ 3.2 An extension to composite pre-change
We justify this assumption by appealing to Huber’s notion of “least favorable distribution pair” [13] in terms of risk, where the risk $R$ for a hypothesis test $\phi$ between $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$ is defined as
$R(P_{0},\phi)=C_{0}P_{0}(\phi\text{ accepts }\mathcal{P}_{1}),\penalty 10000\ \penalty 10000\ R(P_{1},\phi)=C_{1}P_{1}(\phi\text{ accepts }\mathcal{P}_{0}).$ (3.8)
@@
Proposition 3.7.
$(P_{0}^{*},P_{1}^{*})$ is a least favourable distribution pair in terms of risk for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, then Assumption 3.5 holds with $F^{*}=P^{*}$, and $\tau$ being the stopping time of a CUSUM/SR change detector with the likelihood ratio of $P_{1}^{*}$ to $P_{0}^{*}$.
LFD pairs are known to exist for various parametric and nonparametric models, e.g., the monotone likelihood ratio (MLR) families, and Huber’s robust models (including $\epsilon$ neighborhood model, total variation model, and more generally with 2-alternating capacities [14]).
@@
$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}^{*}}\right\}.$ (3.9)
@@
Suppose that the pre-change class $\mathcal{P}_{0}$ and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfy Assumption 3.5 and $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,\forall F_{0}\in\mathcal{P}_{0}$. Then, for any $\alpha\in(0,1),$ $T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, the confidence set $\mathcal{C}$ defined in (3.9), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
@@
We now analyze the size of the confidence set, generalizing Theorem 3.3 to the case where an LFD pair $(F_{0}^{*},F_{1}^{*})$ exists for $\mathcal{P}_{0}$ vs $\mathcal{P}_{1}$. We consider the confidence set in (3.9) with $r_{t}^{*}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}^{*}\geq t)}{N+1}$, $M_{t}$ as defined in (3.4) with $\hat{T}$ as defined in (3.3) using $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})},$ and $S_{n}^{(t)}=\prod_{i=n}^{t-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}$. We redefine $p_{t},\rho_{0}$ and $\rho_{1}$ in terms of the LFD pair: $p_{t}=\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t)$, $\rho_{0}(s)=\mathbb{E}_{F_{0}^{*}}\left(\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\right)^{s}$ and $s_{0}=\arg\min_{s\in(0,1)}\rho_{0}(s)$ and $\rho_{1}(s)=\mathbb{E}_{F_{1}}\left(\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\right)^{s}$ and $s_{1}=\arg\min_{s\in(0,1)}\rho_{1}(s)$. Finally, we call the stopping time $\tau$ well-behaved if it satisfies
$\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ $\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t),\forall t\in\mathbb{N},\forall F_{0}\in\mathcal{P}_{0};$ (3.10a)
$\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},\infty}[\tau\leq t],\forall t,T\in\mathbb{N}\penalty 10000\ \text{ and }$ (3.10b)
$\mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t],\forall t,T\in\mathbb{N},\text{ and }\forall(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ (3.10c)

### v6

@@ 3 A general framework for nonparametric post-change
Let $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ denote the pre- and post-change classes respectively, $\mathcal{P}_{0}\cap\mathcal{P}_{1}=\emptyset$. Our goal is to construct post-detection confidence sets for the changepoint $T$ by inverting tests for $H_{0,t}$, as formalized in Proposition 2.
@@
$M_{t}=\begin{cases}\displaystyle\max\left\{\max_{1\leq j\leq t-1}\prod_{i=j}^{t-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})},\max_{t\leq j\leq\tau}\prod_{i=t}^{j}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right\},&\text{ if }t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (8)
One natural way to extend this test statistic for the general nonparametric setting is to replace the likelihood ratio by e-processes [ramdas2022game, grunwald2024safe], which have emerged as a central tool in anytime-valid inference. While e-processes have been used for sequential testing [wasserman2020universal, ramdas2022testing, saha2024testing], estimation [waudby2023estimating], and change detection [shin2022detectors], we employ them for the first time for changepoint localization. For $t\in\mathbb{N}$, we use two different types of e-processes:
@@
Such powered e-processes for $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ (and vice versa) exist if the total variation distance between the two convex hulls of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ is positive (see e.g., Section 3.4 of [ramdas2024hypothesis]). The likelihood ratio process is the simplest example of an e-process. There is a substantial literature on constructing e-processes in both parametric (e.g., [wasserman2020universal]) and nonparametric settings (e.g., [waudby2023estimating]). For readers less familiar with these developments, we refer to Chapters 6 and 7 of the recent textbook [ramdas2024hypothesis].
With the above ingredients in place, we construct a test statistic for $H_{0,t}$ by comparing it against a competing hypothesis $H_{0,T}$. Suppose $t<T$. Then under $H_{0,t}$ (which is false), the segment $X_{t},\dots,X_{T-1}$ follows a post-change distribution from $\mathcal{P}_{1}$ (because under $H_{0,t}$, the change occurred at $t$), whereas under $H_{0,T}$ (which is true), it follows a pre-change distribution from $\mathcal{P}_{0}$. Outside this segment $X_{t},\dots,X_{T-1}$, both hypotheses agree on the distribution of the observations, and hence those are non-informative for testing between $H_{0,t}$ and $H_{0,T}$. Therefore, we use a forward e-process $R^{(t)}_{T-1}$ that accumulates evidence in favor of $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ over this segment. Similarly, if $t>T$, we use a backward e-process $S^{(t)}_{T}$ to accumulate evidence for $\mathcal{P}_{0}$ against $\mathcal{P}_{1}$, using the segment $X_{T},\dots,X_{t-1}$ . Since the true $T$ is unknown, we consider the maximum over all valid choices of $T$ for both the e-processes, i.e., $\max_{t\leq n\leq\tau}R^{(t)}_{n}$ and $\max_{1\leq n\leq t-1}S^{(t)}_{n}$. We define the final test statistic as the maximum of these two quantities:
$M_{t}=\begin{cases}\displaystyle\max\left\{\max_{t\leq n\leq\tau}R^{(t)}_{n},\max_{1\leq n\leq t-1}S^{(t)}_{n}\right\},&\text{ if }t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (9)
@@
While this bound may be conservative, it is remarkably universal: it holds without any assumptions on the distribution classes $\mathcal{P}_{0}$, $\mathcal{P}_{1}$, or on the detection method $\mathcal{A}$. For constructing a valid confidence set, Proposition 2 implies that we require tests of level $\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t)$, whose value is of course not known to us. To circumvent this issue, we first assume that the pre-change distribution $F_{0}$ is known. This is a common assumption in sequential changepoint literature.
@@ 3.1 Confidence set for $T$ with known pre-change distribution
$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}}\right\}.$ (10)
@@
Theorem 2.
Suppose that the pre-change distribution is known, i.e., $\mathcal{P}_{0}=F_{0}$. Then, for any $\alpha\in(0,1),T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfying $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (10), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{1}\in\mathcal{P}_{1}.$
Now we analyze the size of the confidence set (10) for known pre-change $F_{0}$ and post-change $F_{1}$, with $r_{t}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)}{N+1}$, $M_{t}$ as defined in (8). We provide a bound on the expected size of the confidence set as $N\to\infty$, where $N$ is the number of simulations used to calculate $r_{t}$. Define, $p_{t}=\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$. Note that $r_{t}\to p_{t}$ as $N\to\infty$. Define $\rho_{0}(s)=\mathbb{E}_{F_{0}}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}$ and $\rho_{1}(s)=\mathbb{E}_{F_{1}}\left(\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right)^{s}$. These two quantities measure closeness between $F_{0}$ and $F_{1}$, and thus the problem hardness. Note that for $s=1/2,$ this coincides with Bhattacharyya distance. At $s=0,1$, $\rho_{0}$ and $\rho_{1}$ take value $1.$ Since for $s\in(0,1)$, $h(x)=x^{s}$ is a strictly concave function and since $F_{1}\neq F_{0}$, $\frac{f_{1}(X)}{f_{0}(X)}$ is not almost everywhere constant, we get a strict inequality using Jensen’s: $\rho_{0}(s)=\mathbb{E}_{F_{0}}\left[\left(\frac{f_{1}(X)}{f_{0}(X)}\right)^{s}\right]<\left(\mathbb{E}_{F_{0}}\left[\frac{f_{1}(X)}{f_{0}(X)}\right]\right)^{s}=1.$ Similarly, for $s\in(0,1),\rho_{1}(s)<1$. We now define $s_{0}=\arg\min_{s\in[0,1]}\rho_{0}(s)$ and $s_{1}=\arg\min_{s\in[0,1]}\rho_{1}(s)$, and note that the preceding facts imply that $0<s_{0}<1$ and $0<s_{1}<1$. Our length bound will depend on the expected detection delay, $\Delta_{F_{0},T,F_{1}}:=\mathbb{E}_{F_{0},T,F_{1}}(|\tau-T|\mid\tau\geq T).$ For typical detection algorithms, this quantity often scales like the logarithm of the average run length divided by the KL divergence between $F_{1},F_{0}$. Finally, we call the stopping time $\tau$ as sensitive if
$\penalty 10000\ \mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]\geq\mathbb{P}_{F_{0},\infty}[\tau\leq t],$ (11)
@@
Theorem 3.
In the setting described above, for any change detection method with stopping time $\tau$ satisfying $p_{T}\neq 0$, the size of $\mathcal{C}$} can be bounded as:
$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\cdot\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{F_{0},T,F_{1}}.$ (12)
Additionally, if $\tau$ is sensitive in the sense of (11), the last term in (12) can be improved to $\Psi_{F_{0},T,F_{1}}:=\min\left\{\Delta_{F_{0},T,F_{1}},\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right)\right\}$.
Remark 1.
Suppose $F_{1}=N(\mu_{1},\sigma^{2})$ vs $F_{0}=N(\mu_{0},\sigma^{2})$ and we are using a CUSUM detector with the likelihood ratio (which satisfies the assumption) with average run length $A$. Consider the asymptotic regime when $A\to\infty$, and $T=CA$, for some constant $C>0$. Using the property of the CUSUM detector that $\frac{\tau}{\mathbb{E}_{F_{0},\infty}(\tau)}\stackrel{{\scriptstyle d}}{{\to}}\exp(1)$ as $A\to\infty$, we have $p_{T}=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq CA_{T})\to\exp(-C)$. Also, for the Gaussian distribution, it is easy to verify that $\rho_{0}(s)=\rho_{1}(s)=\exp\left(-s(1-s)K\right)$, where $K=\frac{(\mu_{1}-\mu_{0})^{2}}{2\sigma^{2}}$ and $s_{0}=s_{1}=1/2$. Substituting, we get
@@
To simplify further, consider the difficult regime of small $K$, where $\left(\exp\left(\frac{K}{4}\right)-1\right)\approx K/4$, and $\left(\exp\left(\frac{K}{8}\right)-1\right)\approx K/8$, and take $C=\frac{1}{3}\log(\frac{2}{\alpha})$, so that the above expression is bounded by, ignoring constants, $K^{-1}[\frac{1}{\alpha}+\frac{1}{\alpha^{5/6}}++\frac{1}{\alpha^{1/3}}]$. The fact that the bound grows with $C=T/A$ is empirically verified in Supplementary Section G.1. This behavior is not a limitation of our methodology; rather, it reflects a fundamental difficulty of the problem: $p_{t}$ decays as $C$ increases, and hence, by our duality result (Proposition 2), the tests corresponding to any confidence set will have smaller levels $\alpha p_{t}$, leading to reduced power and, consequently, wider confidence sets.
@@ 3.2 An extension to composite pre-change
We justify this assumption by appealing to Huber’s notion of “least favorable distribution pair” [huber1965robust] in terms of risk, where the risk $R$ for a hypothesis test $\phi$ between $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$ is defined as
$R(P_{0},\phi)=C_{0}P_{0}(\phi\text{ accepts }\mathcal{P}_{1}),\penalty 10000\ \penalty 10000\ R(P_{1},\phi)=C_{1}P_{1}(\phi\text{ accepts }\mathcal{P}_{0}).$ (13)
@@
Proposition 3.
$(P_{0}^{*},P_{1}^{*})$ is a least favourable distribution pair in terms of risk for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, then Assumption 1 holds with $F^{*}=P^{*}$, and $\tau$ being the stopping time of a CUSUM/SR change detector with the likelihood ratio of $P_{1}^{*}$ to $P_{0}^{*}$.
LFD pairs are known to exist for various parametric and nonparametric models, e.g., the monotone likelihood ratio (MLR) families, and Huber’s robust models (including $\epsilon$ neighborhood model, total variation model, and more generally with 2-alternating capacities [huber1973minimax]).
@@
$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}^{*}}\right\}.$ (14)
@@
Suppose that the pre-change class $\mathcal{P}_{0}$ and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfy Assumption 1 and $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,\forall F_{0}\in\mathcal{P}_{0}$. Then, for any $\alpha\in(0,1),$ $T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, the confidence set $\mathcal{C}$ defined in (14), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
@@
We now analyze the size of the confidence set, generalizing Theorem 3 to the case where an LFD pair $(F_{0}^{*},F_{1}^{*})$ exists for $\mathcal{P}_{0}$ vs $\mathcal{P}_{1}$. We consider the confidence set in (14) with $r_{t}^{*}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}^{*}\geq t)}{N+1}$, $M_{t}$ as defined in (9) using $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})},$ and $S_{n}^{(t)}=\prod_{i=n}^{t-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}$. We redefine $p_{t},\rho_{0}$ and $\rho_{1}$ in terms of the LFD pair: $p_{t}=\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t)$, $\rho_{0}(s)=\mathbb{E}_{F_{0}^{*}}\left(\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\right)^{s}$ and $s_{0}=\arg\min_{s\in(0,1)}\rho_{0}(s)$ and $\rho_{1}(s)=\mathbb{E}_{F_{1}}\left(\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\right)^{s}$ and $s_{1}=\arg\min_{s\in(0,1)}\rho_{1}(s)$. Finally, we call the stopping time $\tau$ well-behaved if it satisfies
$\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ $\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t),\forall t\in\mathbb{N},\forall F_{0}\in\mathcal{P}_{0};$ (15a)
$\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},\infty}[\tau\leq t],\forall t,T\in\mathbb{N}\penalty 10000\ \text{ and }$ (15b)
$\mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t],\forall t,T\in\mathbb{N},\text{ and }\forall(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ (15c)